OpenStudy (anonymous):

the circumference of a log is 62.8 in. how long is the side of a cross section of the largest square beam that can be cut from the log.

OpenStudy (anonymous):

sqrt(2) * r what is r ? circumf. = 2*pi*r = 62.8 , find r from here next, why is the side = sqrt(2) * r ? can you tell me?

OpenStudy (anonymous):



OpenStudy (anonymous):

whats the side of the radius ?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):